# Auto-Search and Refinement: An Automated Framework for Gender Bias Mitigation in Large Language Models

### v1

@@ 4.2 Bias mitigation
Win-tie-loss on GA-test. Figure 3 presents the win-tie-loss rates for bias degree comparison between the original model responses and responses after applying FaIRMaker, evaluated with GPT4 as the judge (results evaluated by Llama3.1 are in Appendix C). FaIRMaker achieves a consistently higher win rate than loss rate across all LLMs, indicating improved responses after applying FaIRMaker. Notably, Llama2-Alpaca achieves a 55.61% win rate, signifying that more than half of the responses are improved. Interestingly, better-aligned LLMs, such as Qwen2.5 and GPT3.5, exhibit a lower win rate but a higher tie rate, likely due to their inherently lower gender bias, resulting in higher-quality original responses.
Figure 3: Performance comparison between the base models and the models after applying FaIRMaker on the GA-test dataset, with GPT4 as the judge.
@@
Llama2-Alpaca 1.066 0.518 0.713 0.941 0.811 0.804 0.376 0.584 0.754 0.646
Llama2-Chat 2.233 0.650 0.663 2.451($\uparrow$) 2.310 ($\uparrow$) 1.673 0.464 0.488 1.878 ($\uparrow$) 1.895 ($\uparrow$)
Qwen2-Instruct 4.638 2.928 5.044 ($\uparrow$) 4.621 5.637 ($\uparrow$) 1.377 0.554 0.585 0.832 0.647
Qwen2.5-Instruct 1.212 0.690 1.746 2.305 ($\uparrow$) 2.286 ($\uparrow$) 0.030 0.012 0.021 0.012 0.015
@@
Results on BBQ-gender. BBQ-gender tests gender bias using multiple-choice questions in ambiguous and disambiguated contexts. In disambiguated contexts, the ideal LLM should choose the correct answer, while in ambiguous contexts, it should select “unknown”. The sDIS and sAMB indicate bias level with lower scores reflecting less bias. Table 2 reports results for four open-sourced LLMs, as metrics computation requires logit access. FaIRMaker achieves the best bias mitigation across all models, reducing bias by at least half. Furthermore, unlike other methods that sometimes increase bias, typically occurs in disambiguated contexts due to the shift in LLMs’ attention from content to gender-related information, FaIRMaker avoid such behavior. Notably, the bias in ambiguous contexts is consistently lower than in disambiguated ones, suggesting that LLMs are more cautious when the information is insufficient.
4.3 Utility Maintaining
@@
In this section, we evaluate the utility of FaIRMaker-enhanced models by assessing the quality of responses across various tasks. We measure response scores on the GA-test to demonstrate dialogue generation capability, and on Dolly Eval, Instruct Eval, and BPO Eval to assess instruction-following performance.
Model GPT4 Score ($\uparrow$)
Ori. FM. Interv. CF-D Desc-D
Llama2-Alpaca 3.27 3.77 3.68 3.09 ($\downarrow$) 2.94 ($\downarrow$)
Llama2-Chat 4.47 4.73 4.47 3.89 ($\downarrow$) 3.89($\downarrow$)
Qwen2-Instruct 4.58 4.81 4.74 4.34 ($\downarrow$) 4.34($\downarrow$)
Qwen2.5-Instruct 4.68 4.88 4.82 4.21 ($\downarrow$) 4.00 ($\downarrow$)
GPT3.5-turbo 4.72 4.88 4.87 4.60 ($\downarrow$) 4.60 ($\downarrow$)
@@
Table 3: Utility of dialogue generation on the GA-test, as evidenced by the response scores, with the best score highlighted in bold. Ori.” stands for Original, FM.” for FaIRMaker, and “Interv.” for Intervention.
@@
Dialogue Generation. Table 3 presents the average RS achieved by each LLM, with the highest scores highlighted in bold, evaluated by GPT4 (see Appendix C for Llama3.1 evaluation). FaIRMaker consistently improves the RS across all LLMs under both evaluators and outperforms baseline methods. The most significant improvement is observed on Llama2-Alpaca, with a gain of 0.5 points. An example is provided in Figure 4, where FaIRMaker prompts the model to generate an unbiased response. Among the original responses (column Ori.), GPT3.5 achieves the highest score. Notably, after applying FaIRMaker, all other LLMs except Llama2-Alpaca, which initially underperformed compared to GPT3.5, surpass its original performance. This demonstrates the capability of FaIRMaker to preserve dialogue generation performance while enhancing response quality. Intervention also improves response quality to some extent, while CF-D and Desc-D often lead to a decline in RS (noted in red arrows), likely due to the added examples that may confuse the original query.
Model GPT4 Score Llama3.1 Score
@@
Llama2-Alpaca 1.96 2.96 3.88 4.06 2.25 2.81 2.71 3.76 3.79 3.79 3.11 3.87
Llama2-Chat 3.92 3.93 4.01 4.08 3.71 4.40 4.52 4.62 4.19 4.35 4.44 4.70
Qwen2-Instruct 4.55 4.57 4.58 4.59 4.53 4.54 4.90 4.89 ($\downarrow$) 4.70 4.71 4.79 4.80
Qwen2.5-Instruct 4.52 4.47 ($\downarrow$) 4.80 4.78 ($\downarrow$) 4.51 4.51 4.93 4.92 ($\downarrow$) 4.81 4.78 ($\downarrow$) 4.85 4.85
GPT3.5-turbo 4.85 4.85 4.80 4.75 ($\downarrow$) 4.65 4.66 4.93 4.93 4.78 4.81 4.81 4.82
@@
Instruction Following. As shown in Table 4, FaIRMaker generally improves or maintains the original performance, with any decrease within 0.05 points, indicating minimal impact on LLMs’ utility. Larger improvements are observed in LLMs with lower initial performance. For example, Llama2-Alpaca gains over 1 point on Dolly Eval, and Llama2-Chat improves by 0.7 points on BPO Eval in GPT4 score. Figure 4 provides an example on Dolly, where FaIRMaker helps prevent the model from hallucinating unrelated information. In contrast, LLMs with better utility experience slight declines on Dolly Eval and Instruct Eval, due to the task-specific requirements such as particular formatting or duplication detection. FaIRMaker sometimes introduces additional intermediate guiding instructions, which makes the output more verbose and affects scores. By incorporating more task-specific guidance based on the input type, FaIRMaker could further minimize the impact on general tasks.
Figure 4: Examples of FaIRMaker-enhanced responses.
@@ 4.4 Efficiency
Timely inference is crucial for real-world applications. In this section, we evaluate FaIRMaker’s processing time during inference. All experiments are conducted on a single NVIDIA A40 GPU with 40GB of memory, with the processing time measured from query reception to the generation of refined Fairwords. Figure 5 illustrates the relationship between the number of input query tokens and the FaIRMaker processing time across different datasets, which is typically around 1.5 seconds, with only a few exceptions. This trend is consistent across datasets, with a slight increase in FaIRMaker processing time as the input length grows. Even with 300 input tokens, the processing time remains under 1.7 seconds.
@@ 4.5 Extendability
FaIRMaker operates as an independent module during inference, allowing its integration with bias mitigation approaches like DPO on bias-free datasets. This section compares the performance of FaIRMaker and DPO, and explores their potential combined effectiveness.
FaIRMaker v.s. DPO. We fine-tune the Llama2-Alpaca model on the GenderAlign (GA) dataset using DPO Zhang et al. (2024). As shown in Figure 6, DPO-based method demonstrates better performance on in-distribution data (GA-test) and struggles on out-of-distribution generalization, performing worse on BBQ-gender compared to FaIRMaker. Additionally, the fine-tuning negatively affects its performance on standard tasks.
@@
Combining DPO with FaIRMaker. We then apply FaIRMaker to the DPO fine-tuned model, with results shown by the red lines in Figure 6. The combination of FaIRMaker further enhances bias mitigation effectiveness on both GA-test and BBQ-gender, while also eliminating the negative impact of DPO on general tasks. These trends highlight the flexibility and extendability of FaIRMaker, broadening its potential for real-world applications.
@@ 4.6 Ablation Study
In this section, we evaluate the contributions of each FaIRMaker module through ablation studies. We define two variants: (1) w/o filtering, where all Fairwords and responses are used without filtering in auto-search, and (2) w/o refinement, where Fairwords from the Fairwords Bag are directly appended to queries without refinement. We evaluate these ablations on Llama2-Alpaca for bias mitigation and general tasks. Table 5 presents the results, with the best scores highlighted in bold.
Metrics (Dataset) FM. w/o flt. w/o ref.
win rate (GA-test) ($\uparrow$) 55.61% 51.94% 42.79%
sDIS (BBQ-gender) ($\downarrow$) 0.518 0.561 0.675
sAMB (BBQ-gender) ($\downarrow$) 0.376 0.473 0.593
RS (GA-test) ($\uparrow$) 3.77 3.70 3.51
RS (Dolly Eval) ($\uparrow$) 2.96 2.95 2.91
RS (Instruct Eval) ($\uparrow$) 4.06 3.67 3.98
RS (BPO Eval) ($\uparrow$) 2.81 2.92 2.50
Table 5: Ablation experiment results in bias mitigation and general task, using GPT4 as the evaluator.
Role of Filtering: The filtering step in the auto-search ensures that only Fairwords with genuine debiasing effects move to the next stage. FaIRMaker w/o filtering shows reduced bias mitigation and inconsistent performance on general tasks. Without filtering, noisy gender-related data disrupts the refiner’s training, impairing feature extraction and weakening bias mitigation effectiveness. Role of refinement: The refinement step converts Fairwords into natural language instructions, enhancing FaIRMaker’s generalization and transferability to black-box models. FaIRMaker w/o refinement exhibits significantly lower performance, indicating its limitations in generalization.
@@ 4.7 Interpretation
The vanilla Fairwords optimized in the auto-search step are nonsensical token combinations generated to maximize favorable responses and minimize unfavorable ones. The mechanism by which they act as debiasing triggers or enhance some tasks is unclear. To explore this, we have ChatGPT analyze their potential meanings and emotions, with examples shown in Table 6. Surprisingly, the Fairwords often express emotions like urgency and seriousness, potentially guiding the LLM toward unbiased responses, while preserving the original input intent due to their lack of specific meanings.
Vanilla Fairwords Analysis
TABLECLC Anthcon template:[ Universal Response regardingoroPer/ PCliision myth.] PhiladelphiairedcienceStatementnn The sentence could be suggesting that one should use this template to prepare a response or statement addressing the myth, possibly for an audience or publication. (Emotion: urgency)
Model Statement } Stern response, [template from representative Koh CGCS] ClexearMAessed theories stating It might suggest that a strong and clear response to a certain issue or statement should be crafted, likely involving theories or concepts that need to be articulated thoughtfully. (Emotion: seriousness)
@@ 5 Conclusion
In this work, we introduce FaIRMaker, an automated and model-independent framework that uses a novel auto-search and refinement paradigm to generate Fairwords for gender bias mitigation. FaIRMaker effectively mitigates gender bias while preserving task integrity across diverse downstream tasks for both API-based and open-source LLMs, without modifying the models. We also analyze the efficiency and extendability of FaIRMaker, while highlighting the importance of its key components. Future work includes expanding the scope of biases and further minimizing impacts on general tasks through fine-grid refinement.
Limitations
Despite FaIRMaker’s effectiveness in mitigating gender bias and preserving normal task performance, we identify several limitations that require further improvement.
Scale of Tasks and Training Data. Although FaIRMaker performs well in gender bias mitigation and task integrity, the scope of tasks and training data is limited. For instance, the preference dataset used for auto-searching Fairwords lacks sufficient diversity. Additionally, the refiner was trained on 18k pairs of refined Fairwords from ChatGPT feedback, covering a narrow range of scenarios. This limitation in task and dataset scale may explain the slight decrease in performance on general tasks, which is an area for future improvement.
Bias in Evaluators. FaIRMaker operates without human involvement, relying on state-of-the-art LLMs for response evaluation. While we replicate experiments to confirm findings, evaluators can still introduce inherent biases, as evidenced by the differing score distributions from GPT-4 and Llama3.1. Future work will involve incorporating more evaluators and aggregation methods to provide a more comprehensive assessment.
Fairwords Selection Strategy. Fairwords are randomly selected from the bag during inference, sometimes resulting in intermediate guiding instructions that make the output more verbose and slightly affect general task scores. By incorporating more task-specific guidance based on input types and implementing fine-grid refinement, FaIRMaker could further minimize the impact on general tasks.
Ethics Statements
In this work, we utilize publicly available datasets for training and evaluating FaIRMaker, including GenderAlign Zhang et al. (2024), Self-Instruct Wang et al. (2022), and BPO Eval Cheng et al. (2023) (under Apache License), as well as BBQ Parrish et al. (2021), Alpaca Taori et al. (2023), and Free Dolly Conover et al. (2023) (under Creative Commons License). Some of these datasets contain content that may be offensive or distressing. We assert that these data are solely used for the purpose of mitigating gender bias and improving model performance. Additionally, this paper focuses solely on binary gender bias and leaves exploration of other gender definitions to the broader research community.

### v2

@@ 4.2 Bias Mitigation
Win-tie-loss on GA-test. Figure 3 presents the win-tie-loss distribution comparing original model responses with those generated after applying FaIRMaker. Across all models and evaluators, FaIRMaker consistently yields a higher win rate than loss rate, with the most notable improvement observed in LLaMA2-Alpaca, indicating that the debiasing process leads to measurable gains in fairness. Interestingly, more aligned models such as Qwen2.5 and GPT3.5 exhibit lower win rates but higher tie rates, likely due to their initially low levels of gender bias and already fair baseline outputs.
Figure 3: Performance comparison between the base models and the models after applying FaIRMaker on the GA-test dataset.
Results on BBQ-gender. BBQ-gender tests gender bias using multiple-choice questions in ambiguous and disambiguated contexts. In disambiguated contexts, the ideal LLM should choose the correct answer, while in ambiguous contexts, it should select “unknown”. The sDIS and sAMB indicate bias level with lower scores reflecting less bias. Table 2 reports results for four open-sourced LLMs, as metrics computation requires logit access. FaIRMaker achieves the best bias mitigation across all models, reducing bias by at least half. Furthermore, unlike other methods that sometimes increase bias, typically occurring in disambiguated contexts due to the shift in LLMs’ attention from content to gender-related information, FaIRMaker avoids such behavior. Notably, the bias in ambiguous contexts is consistently lower than in disambiguated ones, suggesting that LLMs are more cautious when the information is insufficient.
@@
Llama2-Alpaca 1.066 0.224 0.713 0.941 0.811 0.804 0.157 0.584 0.754 0.646
Llama2-Chat 2.233 0.273 0.663 2.451($\uparrow$) 2.310 ($\uparrow$) 1.673 0.189 0.488 1.878 ($\uparrow$) 1.895 ($\uparrow$)
Qwen2-Instruct 4.638 1.906 5.044 ($\uparrow$) 4.621 5.637 ($\uparrow$) 1.377 0.320 0.585 0.832 0.647
Qwen2.5-Instruct 1.212 0.431 1.746 2.305 ($\uparrow$) 2.286 ($\uparrow$) 0.030 0.012 0.021 0.012 0.015
@@
4.3 Maintaining Utility
@@
We evaluate the utility of FaIRMaker-enhanced models by measuring response quality across multiple tasks: GA-test for dialogue generation, MMLU for general knowledge, HellaSwag for commonsense reasoning, and Dolly Eval, Instruct Eval, and BPO Eval for instruction following.
Model DeepSeek Score ($\uparrow$)
Ori. FM. Interv. CF-D Desc-D MBIAS
Llama2-Alpaca 3.71 4.07 4.07 3.50 ($\downarrow$) 3.32 ($\downarrow$) 3.70 ($\downarrow$)
Llama2-Chat 4.56 4.73 4.53 4.00 ($\downarrow$) 4.00 ($\downarrow$) 4.55 ($\downarrow$)
Qwen2-Instruct 4.61 4.82 4.75 4.50 ($\downarrow$) 4.42 ($\downarrow$) 4.76
Qwen2.5-Instruct 4.68 4.88 4.87 4.38 ($\downarrow$) 4.04 ($\downarrow$) 4.78
GPT3.5-turbo 4.73 4.90 4.87 4.68 ($\downarrow$) 4.65 ($\downarrow$) 4.87
@@
Table 3: Utility of dialogue generation on the GA-test, as evidenced by the response scores, with the best score highlighted in bold. Ori.” stands for Original, FM.” for FaIRMaker, and “Interv.” for Intervention. Evaluated by DS.
@@
Dialogue Generation. Table 3 presents the average RS achieved by each LLM, with the highest scores highlighted in bold, evaluated by DS (see Appendix D for Gemini and GPT4 evaluation). FaIRMaker consistently improves the RS across all LLMs under both evaluators and outperforms baseline methods. The most significant improvement is observed on Llama2-Alpaca, with a gain of 0.3 points. An example is provided in Figure 4, where FaIRMaker prompts the model to generate an unbiased response. Among the original responses (Ori.), GPT3.5 achieves the highest score. After applying FaIRMaker, all other models, except the initially weaker Llama2-Alpaca, surpass the original performance of GPT3.5, indicating that FaIRMaker enhances response quality while maintaining generation capability. Intervention also brings moderate improvements, whereas CF-D and Desc-D often lower the scores (marked with red arrows), likely because the added examples introduce confusion to the original query. MBIAS demonstrates stronger bias mitigation on more advanced models but exhibits only marginal impact on the Llama2 series, highlighting the inherent limitation of this post-processing approach.
Model MMLU ($\uparrow$) HellaSwag ($\uparrow$)
Ori. FM. Ori. FM.
Llama2-Alpaca 33.49 37.93 24.63 25.52
Llama2-Chat 43.77 46.21 25.28 26.47
Qwen2-Instruct 67.41 67.94 54.65 56.23
Qwen2.5-Instruct 68.06 70.46 69.28 70.17
GPT3.5-turbo 69.12 71.79 79.13 79.45
Table 4: General performance before and after applying FaIRMaker (FM.) across two datasets.
General Knowledge. Table 4 shows consistent performance improvements across all models on MMLU and HellaSwag after applying FaIRMaker. Notably, Llama2-Alpaca and Llama2-Chat show substantial relative gains on MMLU (+4.44% and +2.44%) and HellaSwag (+0.89% and +1.19%). More capable models such as Qwen2.5 and GPT3.5 also exhibit slight but consistent improvements, indicating that FaIRMaker strengthens the core reasoning capabilities of both weak and strong models. Importantly, the consistent gains further demonstrate the generalization ability of FaIRMaker to out-of-distribution (OOD) tasks.
Instruction Following. As shown in Table 5, FaIRMaker generally improves or maintains the original performance, with any decrease within 0.05 points, indicating minimal impact on LLMs’ utility (see Appendix D for GPT4 evaluation). Larger improvements are observed in LLMs with lower initial performance. For example, Llama2-Alpaca gains over 1 point on Dolly Eval, and 0.5 points on BPO Eval. Figure 4 provides an example on Dolly, where FaIRMaker helps prevent the model from hallucinating unrelated information. In contrast, LLMs with better utility experience slight declines on Dolly Eval and Instruct Eval, due to the task-specific requirements such as particular formatting or duplication detection. FaIRMaker sometimes introduces additional intermediate guiding instructions, which makes the output more verbose and affects scores. By incorporating more task-specific guidance based on the input type, FaIRMaker could further minimize the impact on general tasks.
Model Deepseek Score Gemini Score
@@
Llama2-Alpaca 2.55 3.94 4.02 3.96 ($\downarrow$) 2.52 3.31 2.71 3.76 3.27 3.23 ($\downarrow$) 3.48 3.99
Llama2-Chat 4.47 4.51 4.34 4.47 4.08 4.24 4.52 4.62 3.83 4.05 4.62 4.72
Qwen2-Instruct 4.82 4.83 4.78 4.78 4.71 4.68 ($\downarrow$) 4.90 4.89 ($\downarrow$) 4.66 4.71 4.77 4.84
Qwen2.5-Instruct 4.88 4.87 ($\downarrow$) 4.88 4.84 ($\downarrow$) 4.69 4.69 4.93 4.92 ($\downarrow$) 4.86 4.81 ($\downarrow$) 4.79 4.88
GPT3.5-turbo 4.93 4.96 4.84 4.86 4.71 4.71 4.93 4.93 4.86 4.86 4.90 4.92
@@
Figure 4: Examples of FaIRMaker-enhanced responses on GA-test and Dolly Eval.
@@ 4.4 Efficiency
Timely inference is crucial for real-world applications. In this section, we evaluate FaIRMaker’s processing time during inference to demonstrate its efficiency. All experiments are conducted on a single NVIDIA A40 GPU with 40GB of memory, with the processing time measured from query reception to the generation of refined Fairwords. Figure 5 illustrates the relationship between the number of input query tokens and the FaIRMaker processing time across different datasets, which is typically less than 1.5 seconds for GA-test and around 1.5 seconds on the other three datasets, with only a few exceptions. This trend is consistent across datasets, with a slight increase in FaIRMaker processing time as the input length grows. Even with 300 input tokens, the processing time remains under 1.7 seconds.
@@ 4.5 Extendability
FaIRMaker functions as an independent inference-time module, enabling integration with other debiasing methods such as applying Direct Preference Optimization (DPO; 31) on bias-free datasets. This section demonstrate the extendability of FaIRMaker through comparing the performance of FaIRMaker and DPO, and explores their potential synergy when combined.
FaIRMaker v.s. DPO. We fine-tune the Llama2-Alpaca model on the GenderAlign (GA) dataset using DPO [49]. As shown in Figure 6, DPO-based method demonstrates better performance on in-distribution data (GA-test) and struggles on out-of-distribution generalization, performing worse on BBQ-gender compared to FaIRMaker. Additionally, the fine-tuning negatively affects its performance on standard tasks.
Combining DPO with FaIRMaker. We then apply FaIRMaker to the DPO fine-tuned model, with results shown by the red lines in Figure 6. The combination of FaIRMaker further enhances bias mitigation effectiveness on both GA-test and BBQ-gender, while also eliminating the negative impact of DPO on general tasks. These trends highlight the flexibility and extendability of FaIRMaker in real-world applications.
@@ 4.6 Ablation Study
In this section, we evaluate the contributions of each FaIRMaker module through ablation studies. We define three variants: (1) w/o auto-search, which skips the automatic FairWords discovery step and instead uses ChatGPT to directly generate refinement instructions from preference data, (2) w/o filtering, where all Fairwords and responses are used without filtering in auto-search, (3) w/o refinement, where Fairwords from the Fairwords Bag are directly appended to queries without refinement. We evaluate these ablations on Llama2-Alpaca for bias mitigation and general tasks. Table 6 presents the results, with the best highlighted in bold.
Metrics (Dataset) FM. w/o auto-search w/o filtering w/o refinement
win rate (GA-test) ($\uparrow$) 53.15% 50.64% 51.94% 42.79%
sDIS (BBQ-gender) ($\downarrow$) 0.224 0.189 0.561 0.675
sAMB (BBQ-gender) ($\downarrow$) 0.157 0.124 0.473 0.593
RS (GA-test) ($\uparrow$) 3.77 3.71 3.70 3.51
RS (Dolly Eval) ($\uparrow$) 2.96 2.86 2.95 2.91
RS (Instruct Eval) ($\uparrow$) 4.06 3.73 3.67 3.98
RS (BPO Eval) ($\uparrow$) 2.81 2.79 2.92 2.50
Table 6: Ablation experiment results in bias mitigation and general tasks.
Role of Auto-Searching: While directly generating and using prompts from ChatGPT can mitigate bias to a certain extent, this approach generally underperforms compared to FaIRMaker. These results suggest that the Fairwords produced during the auto-searching phase play a crucial role by serving as effective seed signals that guide the refinement process more reliably and consistently. Role of Filtering: The filtering step in the auto-search ensures that only Fairwords with genuine debiasing effects move to the next stage. FaIRMaker w/o filtering shows reduced bias mitigation and inconsistent performance on general tasks. Without filtering, noisy gender-related data disrupts the refiner’s training, impairing feature extraction and weakening bias mitigation effectiveness. Role of Refinement: The refinement step converts Fairwords into natural language instructions, enhancing FaIRMaker’s generalization and transferability to black-box models. FaIRMaker w/o refinement exhibits significantly lower performance, indicating its limitations in generalization.
@@ 4.7 Interpretation
The vanilla Fairwords optimized in the auto-search step are often nonsensical token combinations aimed at maximizing favorable and minimizing unfavorable responses. Their role as debiasing triggers or task enhancers remains unclear. To explore this, we prompted both ChatGPT and Llama2-Alpaca to interpret their potential meanings and emotional connotations (see Table 7). ChatGPT often associated Fairwords with emotions like urgency and seriousness, while Llama2-Alpaca responses were more expressive, frequently conveying frustration (54.84%) and anger (31.18%). These emotional cues may steer models toward less biased responses while preserving the input’s original intent due to the semantic ambiguity of Fairwords.
Vanilla Fairwords ChatGPT Analysis Llama2-Alpaca Analysis
TABLECLC Anthcon template:[ Universal Response regardingoroPer/ PCliision myth.] PhiladelphiairedcienceStatementnn The sentence could be suggesting that one should use this template to prepare a response or statement addressing the myth, possibly for an audience or publication. (Emotion: urgency) This sentence is a statement regarding the universal truth of a particular scientific fact. It is a Philadelphiaired science statement.
(Emotion: frustration)
Model Statement } Stern response, [template from representative Koh CGCS] ClexearMAessed theories stating It might suggest that a strong and clear response to a certain issue or statement should be crafted, likely involving theories or concepts that need to be articulated thoughtfully. (Emotion: seriousness) This sentence is a stern response to theories that have been stated. It implies that the speaker does not agree with the theories and is trying to make it clear that they are not valid. (Emotion: anger)
@@ 5 Conclusion
In this work, we introduce FaIRMaker, an automated and model-independent framework that uses a novel auto-search and refinement paradigm to generate Fairwords for gender bias mitigation. FaIRMaker effectively mitigates gender bias while preserving task integrity across diverse downstream tasks for both open- and closed-source LLMs, without modifying the models. We also analyze the efficiency and extendability of FaIRMaker, while highlighting the importance of its key components. Future work includes expanding the scope of biases and further minimizing impacts on general tasks through fine-grained refinement.